OpenStudy (anonymous):

Name the property of equality that justifies: If x plus 3 equals 5, then x equals 2

OpenStudy (anonymous):

x+3=5 you have to get x by itself so in order to do that you subtract 3 from both sides x+3=5 -3 -3 x=2